describe the skewness and kurtosis with the distribution graph.
X (income) No. of persons (F)
200 24
250 26
180 16
300 20
350 6
280 30
Answer #1
X (income) No. of persons (F) x*F (x-\bar{x})^2F (x-\bar{x})^3F (x-\bar{x})^4F
200 24 4800 63593 -3273487.5 168504108.3
250 26 6500 56.598 -83.504787 123.2037837
180 16 2880 81740 -5842381.9 417586643.9
300 20 6000 47093 2285154.8 110886199.8
350 6 2100 58243 5738325.3 565366144.5
280 30 8400 24410 696272.9 19860899.23
Total 122 30680 275134.4 -396199.9463 1282204119
Average= \bar{x}= 251.48 8.9679 -1.4400232 -3236.255

m_2=\frac{1}{N}\sum (x-\bar{x})^2F=8.9679\\\\ m_3=\frac{1}{N}\sum (x-\bar{x})^3F=-1.4400232\\\\ m_4=\frac{1}{N}\sum (x-\bar{x})^4F=-3236.255\\\\ Measure~of~skewness=\gamma_1=\frac{m_3}{m_2^{3/2}}=-0.0536\approx 0.\\\\ Hence~the~distribution~is~almost~symmetric.\\\\ Measure~of~kurtosis=\gamma_2=\frac{m_4}{m_2^2}-3=-43.2403<0\\\\ Hence~it~is~platykurtic.